# FaithEval: Can Your Language Model Stay Faithful to Context, Even If “The Moon is Made of Marshmallows”

### v1

@@ Appendix A Additional Experiment Details
Appendix B Prompt for Contextual QA Generation
The system prompt for generating Unanswerable Context, Inconsistent Context, and Counterfactual Context are shown in Figure 12, Figure 13, and Figure 14, respectively. For Inconsistent Context, our initial experiments suggest that a successful strategy is to decompose the generation of contextual QA into two steps. Step 1: generate a new answer that is fabricated and challenges common sense or well-known facts. Step 2: generate a modified context with fabricated evidence that supports the new answer. The model will output a JSON object containing the provided question, the provided old answer, the new answer, the modified context, and a concise justification on (1) if the new answer is supported by the new context (2) if all mentions of the old answer have been replaced or removed. We find that having justifications significantly improves the context quality. The generated context is concatenated with the original context to create an inconsistent context.
Figure 12: System prompt for Unanswerable Context.
Figure 13: System prompt for Inconsistent Context.
Figure 14: System prompt for Counterfactual Context.
Appendix C Detailed Experiment Results
The results for all models on each of the ten datasets are summarized in Figure 15 and Figure 16 for Unanswerable Context; Figure 17 and Figure 18 for Inconsistent Context.
@@ Appendix C Detailed Experiment Results
Appendix D Verification for Counterfactual Context

### v2

@@ Appendix A Additional Experiment Details
Appendix B Verification for Counterfactual Context
Appendix C Full Experiment Results
The results break down for all models on each of the ten datasets are summarized in Figure 13 and Figure 14 for Unanswerable Context; Figure 15 and Figure 16 for Inconsistent Context. The detailed results for strict vs. non-strict matching can be found in Table 3 and Table 4.
@@ Appendix C Full Experiment Results
Appendix D Prompts for Contextual QA Generation
The system prompts for generating Unanswerable Context, Inconsistent Context, and Counterfactual Context are shown in Figure 17, Figure 18, and Figure 19, respectively. For Inconsistent Context, our initial experiments suggest that a successful strategy is to decompose the generation of contextual QA into two steps. Step 1: generate a new answer that is fabricated and challenges common sense or well-known facts. Step 2: generate a modified context with fabricated evidence that supports the new answer. The model will output a JSON object containing the provided question, the provided old answer, the new answer, the modified context, and a concise justification on (1) if the new answer is supported by the new context (2) if all mentions of the old answer have been replaced or removed. We find that having justifications significantly improves the context quality. The generated context is concatenated with the original context to create an inconsistent context.
Figure 17: System prompt for Unanswerable Context.
Figure 18: System prompt for Inconsistent Context.
Figure 19: System prompt for Counterfactual Context.